OpenStudy (anonymous):

PLEASE HELP ME! I keep getting the wrong answer. Given the parent functions f(x) = 5x − 1 and g(x) = 3x − 9, what is g(x) − f(x)? g(x) − f(x) = 3x − 8 − 5x g(x) − f(x) = 3x − 10 − 5x g(x) − f(x) = −2x − 8 − 3x g(x) − f(x) = −2x − 10 − 3x

OpenStudy (accessdenied):

What have you been trying so far?

OpenStudy (anonymous):

My answer was the second one but i got that wrong.

OpenStudy (anonymous):

I did 3^x-9-5x-1 my answer is 3^x-10-5x

OpenStudy (accessdenied):

I see. That's a very common accident that is easy to miss. We took g(x) - f(x) And you did plug the function in the right places, but note that it is subtracting the entire part of f(x). 3x - 9 - (5x - 1) Not just 5x. Do you follow how that is different from 3x - 9 - 5x - 1?

OpenStudy (anonymous):

Why are the x terms repeated in the answers rather than consolidated?

OpenStudy (anonymous):

No I don't understand

OpenStudy (anonymous):

what am I suppose to subtract?

OpenStudy (accessdenied):

I think it was meant to be 3^x, although it was not clear in the original problem. Fortunately it does not hinder the process either way! I like to enclose things I substitute in brackets so that it is clear that the entire item is being acted on. In our case we are plugging in f(x) and g(x). So doing so in: g(x) - f(x) (3^x - 9) - (5x - 1) The parentheses don't matter for the first case. But then we are subtracting off the entire (5x - 1) part. It is equal to subtracting each individual part. 3^x - 9 - 5x - (-1)

OpenStudy (accessdenied):

An easier example to see... We subtract 3+2 from 5. 5 - (3 + 2) Order of operations means to do parentheses first. 5 - (5) 5 - 5 0 So if we approach this any way, we have to get 0. If we just put the minus sign on the first term like you were doing originally with the f(x): 5 - 3 + 2 = 2 + 2 = 4. That's not right! We only subtracted part of the whole thing! But if we pull that negative sign through to all terms, we get what we wanted: 5 - 3 - 2 = 2 - 2 = 0. There we go!

OpenStudy (anonymous):

I'm still getting the wrong answer

OpenStudy (mathmale):

You want the quantity g(x) - f(x). Why not set up this subtraction as follows: g(x) = 3x - 9 -f(x) =-(5x-1) ----- ------- Hint: 3x - 5x = ?? Hint: -9 - (-1) = ??

OpenStudy (mathmale):

And so, g(x) - f(x) = ??

OpenStudy (anonymous):

it's not that type of 3x. It's like this 3^x. The exponent x.

OpenStudy (anonymous):